# SAGe: A Lightweight Algorithm-Architecture Co-Design for Alleviating Data Preparation Overheads in Large-Scale Genome Analysis

### v1

@@ 5.2. SAGe’s Hardware Design
Fig. 10 shows the structure of SAGe’s hardware accelerator, consisting of three key components. 1 Scan Unit (SU) sequentially scans through position arrays to find the mismatch information. To do so, it scans through the position guide array and as soon as it reaches a zero bit, it detects that a new element of the guide array is read. Based on this element’s value, SU determines how many bits to read from the position array (e.g., as shown in the example of Fig. 7). The SU then adds the newly-read value of the position array to the previously decoded mismatch position to find the new mismatch position. 2 Read Construction Unit (RCU) receives the position information from SU and reconstructs the full reads by plugging the mismatches into the correct positions of the consensus sequence. RCU sequentially scans through the consensus sequence and whenever it reaches a mismatch position, it replaces the base in that position with the mismatching base. 3 Control Unit (CU) coordinates operations between SU and RCU.
Due to the streaming access patterns of SU and RCU operations, they only require small registers to buffer a small part of each array. The Guide Array Register and Array Registers each consist of eight bits since that is the largest element size used in the arrays in our compression scheme design. The Configuration Register is also eight bits since we load the configuration parameters in eight-bit chunks. We consider a Consensus Sequence Register of size 150 base pairs since this is the largest read size in most short read sequencing datasets (glenn2011field, ; quail2012tale, ). For short reads longer than this size and for long reads, we reconstruct the reads in 150-base-pair chunks.
@@
When receiving a request to access a read set, SAGe’s per-channel accelerators start streaming data from flash channels. First, SU reads the configuration parameters of the arrays and guide arrays ( 1 in Fig. 10). This includes the encoded associations between arrays and guide arrays (e.g., between MPGA and MPA, as shown in Fig. 7). Second, SU scans through the position arrays ( 2) as detailed in §5.2, retrieves the matching positions and mismatch positions ( 3), and sends them to RCU ( 4). Meanwhile, RCU scans through the consensus sequence and the base/type array ( 5). Whenever RCU detects an indel (i.e., when the mismatch base in MBTA equals the consensus sequence base at the corresponding mismatch position, as detailed in §5.1.2), it sends a signal to SU to decode information about the indel and its length accordingly 6. Finally, based on all the mismatch information (i.e., mismatch position, type, and base), RCU reconstructs each read by substituting or inserting mismatches into the correct positions in the consensus sequence, or by deleting the relevant bases if the mismatch is a deletion ( 7). Based on the interface command (§5.3) from the genome analysis system, RCU flexibly constructs the reads in the desired format and sends them for genome analysis ( 8).
@@ 5.3. Interface Commands
SAGe_Write: It is a specialized write command that is used when writing genomic data to the SSD and updates SAGe FTL’s mapping metadata.
@@
5.4. SAGe’s Flash Translation Layer
SAGe FTL needs simple changes to the conventional FTL for handling genomic data and does not interfere with the baseline vendor-specific features of the FTL when handling non-genomic data. SAGe FTL dedicates each block in the SSD to either genomic data or non-genomic data. The SSD detects genomic data is being accessed when it receives a request via SAGe’s interface commands (§5.3). For all other types of data, SAGe does not impact the functionality of the baseline FTL and controller, and the SSD can be accessed similarly to a conventional general-purpose SSD.
Data Placement. SAGe performs data placement for genomic data to leverage the maximum throughput of the SSD and the accelerators of SAGe. To this end, when writing a compressed genomic read set, SAGe uniformly partitions the data across the flash channels in a round-robin manner. This efficient partitioning is enabled by SAGe’s sequential access pattern to all of its data structures. Each partition of the consensus sequence, along with the compressed mismatch information of the reads mapped to that partition, is then placed in a separate channel. Fig. 11 shows an example of how compressed genomic datasets are organized in blocks across SSD channels, dies, and planes. When storing a compressed read set across all channels, SAGe FTL writes data in a round-robin manner such that the active block (i.e., blocks available for writes) in different channels have the same page offset. This placement then enables SAGe FTL to simply perform efficient multi-plane read operations when accessing data across all channels and leverage the SSD’s full bandwidth. With this efficient data placement, and given that each genomic read in a genomic dataset can be analyzed independently, SAGe concurrently streams through data from each channel and performs decompression and reformatting (§5.2).
Figure 11. Data layout in SAGe.
SSD Management Tasks. When accessing and preparing genomic data, SAGe does not write any new data and only accesses the read set sequentially. Therefore, during this phase, SAGe does not need to perform write-related management (e.g., garbage collection (kim2020evanesco, ; cai2017error, ; park-dac-2016, ; park-dac-2019, ) and wear-leveling (chang2007efficient, ; cai-insidessd-2018, )). SAGe’s accelerators are located in the SSD controller and access data after ECC decoding.
When performing garbage collection on genomic data, we must select victim blocks such that we preserve SAGe’s capabilities for efficient multi-plane operations. This requires maintaining the same page offset across all blocks in a parallel unit. Given that genomic read sets do not require partial updates, and that they are accessed sequentially, it is straightforward to realize this efficient garbage collection. SAGe FTL performs garbage collection in a grouped manner and selects every block in the parallel unit as a group of victim blocks, which are then sequentially rewritten. SAGe selects this group of victim blocks in the order they were originally written, as indicated by their logical address sequence in the dataset.
SAGe performs other management tasks needed for ensuring reliability (e.g., refresh to prevent uncorrectable errors (luo2018improving, ; luo-hpca-2018, ; cai2013error, ; cai2012flash, ; cai2017error, ; ha2015integrated, )) before or after its operations during the access of a read set (i.e., decompressing and preparing reads in the desired format). This is because (i) the duration of each SAGe process when accessing a read set, even if the read set occupies entire the SSD capacity, is significantly smaller than the SSD manufacturer-specified threshold for reliable retention age (e.g., one year (micron3dnandflyer, )), and because (ii) SAGe avoids read disturbance errors (cai2015read, ) by accessing a read set sequentially and with low reuse. By avoiding these management tasks during SAGe’s operations when preparing genomic data and leveraging SAGe’s efficient data placement, SAGe’s hardware accelerators can effectively leverage the SSD’s full bandwidth.
5.5. SAGe with Distributed Storage Systems
Due to SAGe’s efficient streaming access patterns and the ability to analyze different batches of genomic reads independently, SAGe is well-suited for integration with a distributed system. We elaborate on this by discussing three key aspects. The first aspect is the use of multiple SSDs. Since all data structures in SAGe are accessed sequentially, they can be evenly partitioned between multiple SSDs. This enables the efficient utilization of multiple SSDs in the system. The second aspect is inter-node communication. Since each batch of genomic reads can be analyzed independently, computation on the batches of data stored in each storage device is independent of data in other devices. This eliminates the need for frequent communication and synchronization between different nodes. The third aspect is accessing the storage system. From the perspective of the compute node, using distributed storage is the same as local storage (other than the latency/bandwidth properties across configurations). Therefore, SAGe can leverage the same type of interface command provisions (§5.3) for distributed storage systems as well.
@@ 6. Methodology
Evaluated Systems. We show the benefits of SAGe on end-to-end performance (preparing the data and performing read mapping on it) when it is integrated with a state-of-the-art read mapping accelerator, GEM (chen2023gem, ). For data preparation, we evaluate configurations with the following decompression tools: (ipigz: A parallel version (adler2015pigz, ) of gzip, a commonly-used general-purpose compressor; (ii(N)Spr: Spring (chandak2018spring, ) and NanoSpring (Meng2023, ), state-of-the-art compressors for short and long reads, respectively; (iii(N)SprAC: (N)Spr integrated with a BWT-accelerator. There are various accelerators available for BWT (e.g., (qiao2019fpga, ; guo2013gpu, ; zhao2017streaming, )). In our experiment, we consider an idealized accelerator that can fully eliminate the BWT execution time from (N)Spr; (iv0TimeDec: an ideal decompression tool (with zero time needed for decompression, and with a compression ratio similar to the state-of-the-art genomic compressor) outside the SSD, to show the limitations of decompression tools that cannot be efficiently implemented inside the SSD, regardless of their optimized performance outside the SSD; [^13] As a strongly conservative evaluation, 0TimeDec can also serve as an idealized representation of the closed-source industry software, called ORA (dragenora, ), recently developed for Illumina (illumina, ) short read (de)compression. ORA works similarly to other genomic compressors (as discussed in §2.3), but uses a different backend compressor for mismatch information, which still suffers from limitations for in-storage implementation (detailed in §4). We cannot directly use ORA as our short-read decompression baseline because it decompresses both DNA bases and quality scores, whereas SAGe, like many other tools, targets only the bases (as discussed in §5.1.5). Since ORA’s source code and implementation details are unavailable, modifying it to target only bases is infeasible, which prevents a fair quantitative comparison against it. (vSGSW: SAGe with its decompression implemented in software and running on the host system, to show the benefits of SAGe’s algorithmic optimizations; and (viSG: SAGe’s full implementation, with its decompression implemented in hardware. In all experiments except Fig. 13, SAGe’s hardware accelerator is inside the SSD. In Fig. 13, we compare both SAGe configurations (i.e., with its accelerator inside and outside the SSD) to evaluate their benefits on systems with different SSD bandwidth configurations. [^14] For fair comparisons, in all SAGe implementations, we use the same consensus sequence as (N)Spr. SAGe can flexibly work with any consensus sequence.
We demonstrate the benefits of SAGe’s in-storage implementation by evaluating its integration with a state-of-the-art ISP system for genomics, GenStore (mansouri2022genstore, ). GenStore prunes reads that do not require the costly read mapping process using low-cost and accurate in-storage filters, providing two key benefits: (i) mitigating data-movement overheads and (ii) reducing the computation burden from the rest of the system.
@@
Datasets. We use real-world genomic data, with both short and long read sets. Storage systems must handle numerous read sets. Therefore, compressing each read set is essential to reduce the overall storage burden, regardless of their individual sizes. Thus, we analyze read sets of varying sizes, as listed in Table 3.
Performance. We design a simulator that models all of SAGe’s components, including accessing NAND flash chips, in-storage accelerators (both for SAGe’s hardware units and the in-storage filters (mansouri2022genstore, )), the SSD’s internal DRAM, and host-SSD and accelerator-SSD interfaces. Using this methodology, as also demonstrated in prior works (e.g., (mansouri2022genstore, ; park2022flash, ; megis, )), enables us to incorporate state-of-the-art system configurations in our analysis. We feed the latency and throughput of each component to this simulator. For the components in the hardware-based operations, we implement SAGe’s logic components in Verilog and synthesize them using the Synopsys Design Compiler (synopsysdc, ) with a 22 nm library (22gf, ). We use two state-of-the-art simulators, Ramulator (kim2016ramulator, ; ramulatorsource, ) to model the SSD’s internal DRAM, and MQSim (tavakkol2018mqsim, ; mqsimsource, ) to model the SSD’s internal operations. For the hardware read mapper, we use the throughput numbers reported by the original paper (chen2023gem, ). For the components in the software-based operations (software decompression), we measure performance on a real system, an ${}^{\text{\textregistered}}$ ${}^{\text{\textregistered}}$ 7742 CPU (amdepyc, ) with 128 physical cores and 1.5-TB DRAM. We measure the performance of the software tools on this high-end system using their best-performing thread counts.
SSD Configurations. In our real system experiments, we use a state-of-the-art SSD (samsungPM1735, ) with PCIe Gen4 (PCIE4, ). In Fig. 13, to analyze SAGe’s benefits in systems with SSDs of different interface bandwidths, we also use an SSD (samsung870evo, ) with a SATA3 interface (SATA, ). For the MQSim simulations of the in-storage tasks, we faithfully model the SSD configurations based on the configurations in Table 1.
@@ 7.1. Performance
End-to-End Application Performance. Fig. 12 shows the end-to-end mapping performance (i.e., preparing data and performing read mapping on it). All configurations are integrated with a read mapping accelerator outside the SSD. For SAGe, we also evaluate an additional configuration in which it also integrates with a genomic in-storage filter (ISF). [^15] We use GenStore-EM for filtering exactly-matching short reads (mansouri2022genstore, ). We use GenStore-NM for filtering non-matching long reads in a contamination search read mapping use case presented in the original work (mansouri2022genstore, ). In this case, SAGe sends the unfiltered reads (i.e., reads that could not be pruned by ISF) in a decompressed 2-bit encoded format [^16] SAGe sends reads containing N in a 3-bit encoded format (§5.1.4). to the mapper outside the SSD. Speedup is normalized to the genomics-specific baseline (N)Spr.
Figure 12. End-to-end speedup for different read sets.
@@
We make seven key observations. First, SAGe provides significant performance benefits compared to both general-purpose and genomics-specific baselines. SG leads to 12.3$\times$ and 3.9$\times$ average speedup over pigz and (N)Spr, respectively. Second, SG leads to 3.0$\times$ average speedup over (N)SprAC, while requiring only lightweight hardware and efficient access patterns. Third, by efficiently integrating with the in-storage filter, SG+ISF leads to 30.7$\times$, 9.9$\times$, and 7.5$\times$ average speedups compared to pigz, (N)Spr, and (N)SprAC, respectively. Fourth, due to its efficient sequential access patterns, SAGe’s implementation in software also leads to large speedups. On average, SGSW outperforms (N)Spr by 2.4$\times$. Fifth, despite its benefits, SGSW suffers from two key limitations: (i) decompression still significantly impacts end-to-end performance. SGSW leads to 1.6$\times$ average (up to 3.8$\times$) slowdown compared to the configuration without decompression overhead (0TimeDec), and (ii) by not executing inside the SSD, it does not enable integration with genomic ISP designs. SGSW performs, on average, 4.1$\times$ (up to 7.8$\times$) worse than SG+ISF. Sixth, SAGe can fully hide the decompression time in the end-to-end execution pipeline since the decompression time on a batch of reads in a read set is smaller than the mapping time for the previous batch. Note that, as explained in §3, I/O accesses, decompression, and read mapping work in a pipelined manner. Therefore, the end-to-end throughput is determined based on the slowest stage. 0TimeDec and SG achieving the same performance shows that in both cases, decompression is no longer the slowest stage of the end-to-end pipeline. The hardware implementation of SAGe decompression provides greater speedup than its software by handling bitwise operations (on SAGe’s arrays and guide arrays) more efficiently and managing data flow more effectively in a fine-grained manner. Seventh, regardless of how much decompression tools are optimized for performance outside the SSD, if they are unsuitable for in-SSD adoption (as discussed in §4), they miss out on the benefits of integration with genomic ISP systems. By efficiently integrating with ISF, SG+ISF achieves 2.5$\times$ larger average speedup compared to 0TimeDec.
Figure 13. Ablation analysis of SAGe’s optimizations.
Ablation Study. We perform an ablation study on SAGe’s optimizations by evaluating different configurations: (iSGSW: SAGe purely in software, i.e., only with its algorithmic optimizations, (iiSGout: SAGe with its hardware units, outside the SSD, (iiiSGin: SAGe with its hardware units, inside the SSD, (ivSGin+ISF: SGin integrated with the in-storage filter. [^17] As detailed in §6, we use GenStore (mansouri2022genstore, ) as the in-storage filter. Fig. 13 shows the end-to-end read mapping performance (i.e., preparing data with SAGe and performing read mapping on it). All configurations are integrated with a read mapper accelerator outside the SSD. Note that SAGe’s configurations outside the SSD only integrate with the outside-SSD mapper since their data preparation happens outside the SSD, preventing them from integration with ISF. We evaluate all configurations with a PCIe Gen4 SSD and a cost-optimized SATA SSD, which has significantly lower interface bandwidth than the PCIe SSD. Speedup is normalized to (N)Spr. We make four observations. First, due to its efficient access patterns, SGSW leads to large benefits compared to (N)Spr, providing 2.4$\times$ speedup. Second, SGout provides additional large speedups of up to 3.8$\times$ over SGSW. Third, for the case with low SSD bandwidth (i.e., the SATA SSD), SGout leads to larger benefits than SGin. Fourth, when integrated with ISF, SGin provides large benefits. SGin+ISF leads to better performance than SGout in all cases (providing up to 10.1$\times$ speedup), except when both of the following conditions are met: (i) the input and application do not largely take advantage of in-storage processing (i.e., in this case, ISF does not filter many reads in the read set), and (iithe SSD’s limited external bandwidth bottlenecks performance (e.g., RS1 and RS4 with the SATA SSD). In these cases, the SGout configuration should be used for better performance by decompressing data outside the SSD, thus avoiding moving larger decompressed data through the limited-bandwidth interface.
@@
Performance with Multiple SSDs. Since all data structures in SAGe are accessed sequentially, they can be disjointly partitioned between multiple SSDs. Fig. 14 shows SAGe’s end-to-end performance with multiple SSDs, normalized to (N)Spr. First, SAGe maintains its large speedup when data is partitioned among multiple SSDs. Second, SAGe’s benefits improve for some datasets (RS3 and RS5) since more SSDs improve the performance of ISF, which is at the critical path of their end-to-end performance.
@@
Performance with Distributed Storage. Due to SAGe’s streaming accesses, the key factor of a distributed storage system impacting performance is its bandwidth. We consider two different configurations: (i) a high-end Lustre-based (lustre, ) distributed storage system connected to an InfiniBand network (infiniband, ), achieving a 10 GB/s bandwidth, and (ii) a low-end distributed storage system connected to an Ethernet network (10gbe_ethernet, ), achieving a 10 Gbps bandwidth. Fig. 15 shows the speedup of different SAGe configurations over (N)Spr. We observe that based on the available bandwidth, SGin or SGout can be selected to provide large benefits. For cases with limited bandwidth and with inputs that do not largely take advantage of in-storage processing (e.g., RS1 and RS4 with Ethernet-attached system), SGout should be used for better performance (2.6$\times$ average speedup over (N)Spr) by decompressing data outside the SSD. In all other cases, SGin should be selected to enable efficient integration with ISF, leading to 9.19$\times$ average speedup over (N)Spr.
Figure 15. Speedup with distributed storage configuration.

### v2

@@ 5.2. Hardware for Data Preparation
Fig. 10 shows the structure of SAGe’s hardware accelerator, consisting of three key components. 1 Scan Unit (SU) sequentially scans through position arrays and guide arrays to find the mismatch information. 2 Read Construction Unit (RCU) receives the mismatch information from SU and reconstructs the full reads by plugging the mismatches into the correct positions of the consensus. 3 Control Unit (CU) coordinates operations between SU and RCU. Due to the streaming access patterns of SU and RCU, they only require small registers to buffer a small part of each array. The Guide Array Register and Array Registers each consist of eight bits since that is the largest element size used in the arrays in our compression scheme. The Configuration Register is also eight bits since we load the configuration parameters in eight-bit chunks. We consider a Consensus Sequence Register of size 150 base pairs since this is the largest read size in most short read sequencing datasets (glenn2011field, ; quail2012tale, ). For short reads longer than this and for long reads, we reconstruct the reads in 150-base-pair chunks.
@@
When receiving a request to access a genomic read set, SAGe’s hardware components start streaming data from flash channels. First, SU reads the configuration parameters of the arrays and guide arrays ( 1 in Fig. 10). This includes the encoded associations between arrays and guide arrays (e.g., between MPGA and MPA, as shown in Fig. 7). Second, SU scans through the position arrays ( 2), retrieves the matching positions and mismatch positions ( 3), as detailed in §5.1, and sends them to RCU ( 4). Meanwhile, RCU scans through the consensus sequence and the base/type array ( 5). Whenever RCU detects an indel (i.e., when the mismatch base in MBTA equals the consensus sequence base at the corresponding mismatch position, as detailed in §5.1.2), it sends a signal to SU to decode information about the indel and its length accordingly 6. Finally, based on all the mismatch information (i.e., mismatch position, type, and base), RCU reconstructs each read by substituting or inserting mismatches into the correct positions in the consensus sequence, or by deleting the relevant bases if the mismatch is a deletion ( 7). Based on the interface command (§5.4) from the genome analysis system, RCU flexibly constructs the reads in the desired format and sends them for genome analysis ( 8).
5.3. Data Layout
SAGe requires simple changes to the baseline flash translation layer (FTL) [^10] We consider NAND flash-based SSDs, which are the common technology in modern storage systems. We expect that SAGe’s ideas would also provide benefits with other storage technologies. to enable efficient data layout and leverage the storage system’s full bandwidth. SAGe FTL dedicates each block in the SSD to either genomic or non-genomic data. The SSD detects that genomic data is being accessed when it receives a request via SAGe’s interface commands (§5.4). For all other types of data, SAGe does not interfere with the baseline vendor-specific features of the FTL and, thus, can be accessed similarly to a conventional SSD.
Figure 11. Data layout in SAGe.
When writing a compressed genomic read set, SAGe uniformly partitions data across the SSD’s flash channels. This efficient partitioning is enabled by SAGe’s sequential access pattern to all of its data structures. Each partition of the consensus, along with the compressed mismatch information of the reads mapped to that partition, is placed in a separate channel. Fig. 11 shows an example of how compressed genomic datasets are organized in blocks across SSD channels, dies, and planes. When storing a compressed read set, SAGe FTL writes data in a round-robin manner such that the active block (i.e., blocks available for writes) in different channels have the same page offset. This placement enables SAGe to simply perform efficient multi-plane read operations when accessing data across all channels and leverage the SSD’s full bandwidth.
Garbage Collection. When performing garbage collection on genomic data, we must select victim blocks such that we preserve SAGe’s capabilities for multi-plane operations. This requires maintaining the same page offset across all blocks in a parallel unit. Given that genomic read sets do not require partial updates, and that they are accessed sequentially, it is straightforward to realize this efficient garbage collection. We perform garbage collection in a grouped manner and select every block in the parallel unit as a group of victim blocks, which are then sequentially rewritten. SAGe selects this group of victim blocks in the order they were originally written, as indicated by their logical address sequence.
Other Management Tasks. In SAGe’s design, typical SSD management tasks do not reduce SSD bandwidth during genomic data access. When accessing and preparing genomic data, SAGe does not write any new data and only accesses the read set sequentially. Therefore, during this phase, SAGe does not need to perform write-related management (e.g., garbage collection (kim2020evanesco, ; cai2017error, ; park-dac-2016, ; park-dac-2019, ) and wear-leveling (chang2007efficient, ; cai-insidessd-2018, )). SAGe performs other management tasks needed for ensuring reliability (e.g., refresh to prevent uncorrectable errors (luo2018improving, ; luo-hpca-2018, ; cai2013error, ; cai2012flash, ; cai2017error, ; ha2015integrated, )) before or after its operations during the access of a read set. This is because (i) the duration of each SAGe process when accessing a read set, even if the read set occupies the entire SSD capacity, is significantly smaller than the SSD manufacturer-specified threshold for reliable retention age (e.g., one year (micron3dnandflyer, )), and because (ii) SAGe avoids read disturbance errors (cai2015read, ) by accessing a read set sequentially and with low reuse.
@@ 5.4. Interface Commands
SAGe_Write: It is a specialized write command for writing genomic data to the SSD and updates SAGe FTL’s mapping metadata.
@@
6. Case Studies of SAGe’s Hardware Integration with Genome Analysis Systems
Figure 12. SAGe’s integration with genome analysis systems.
SAGe can be seamlessly integrated with a broad range of genome analysis systems. Fig. 12 shows three examples. In the first case ( 1 in Fig. 12), SAGe’s hardware units are connected to the genome analysis system via a PCIe interface (i.e., similar to how individual accelerators or GPUs connect to the system). In the second case ( 2), SAGe’s hardware is integrated on the same chip as the genome analysis accelerator (which can be a stand-alone ASIC, or an FPGA, etc.), allowing for tighter integration and saving PCIe slots, which are scarce and can be contested by other system components (e.g., SSDs, GPUs, FPGAs, or other accelerators). This integration is easily possible due to SAGe’s low hardware costs. For example, SAGe’s hardware units consume only 2.5% of the lookup tables and 0.8% of the flip-flops of a mid-range FPGA (fpgamidrange, ). In both cases, SAGe does not impose additional burden on the memory system, which is a critical resource for genome analysis tasks. As data pages transfer from SSD to memory (either the host’s DRAM or the accelerator’s memory), SAGe can decompress data with streaming accesses and, thus, not rely on (i) buffering large amounts of data, or (ii) performing many memory accesses to large structures due to random look-ups for matching patterns.
SAGe’s lightweight design enables it to easily integrate with genome analysis systems, even in resource-constrained environments. As an example, in 3, we demonstrate SAGe’s integration with an in-storage NDP genome analysis system. In this example, SAGe’s hardware components are located on the SSD controller, integrated with the genome analysis system, and access data after ECC. We adopt two optimizations to fetch data to SAGe’s hardware such that SAGe (i) leverages the SSD’s full bandwidth and (ii) does not require costly resources in the SSD’s constrained environment. First, similar to (zou2022assasin, ), SAGe’s hardware operates on data fetched from NAND flash chips without needing to buffer them in the SSD’s low-bandwidth, single-channel, internal DRAM. Second, as we no longer rely on DRAM buffers, we must minimize the SRAM buffer sizes necessary for performing computations on NAND flash data streams. We do so based on a key feature enabled by SAGe: SAGe’s operations are lightweight and perform only streaming accesses, eliminating the need to buffer large amounts of data. Therefore, SAGe performs operations on flash data streams using a lightweight double-buffering technique. We use one register to load a chunk of data, and another to hold the subsequent chunk being read from the flash chips. We use two 64-bit registers as they are sufficient to fully utilize the per-channel hardware components.
@@ 7. Methodology
Evaluated Systems. We show the benefits of SAGe on the end-to-end performance (i.e., for preparing data and performing genome analysis on it). For genome analysis, we consider two configurations: (i) GEM (chen2023gem, ): a state-of-the-art read mapping accelerator, (ii) GenStore (mansouri2022genstore, ): a state-of-the-art NDP genome analysis accelerator in a resource-constrained environment, which prunes reads that do not require expensive computation directly inside the SSD, sending only the remaining reads to the read mapper (i.e., GEM) in a decompressed 2-bit encoded format. [^11] SAGe sends reads containing N in a 3-bit encoded format (§5.1.4).
@@
For data preparation, we use the following decompression tools: (ipigz: A parallel version (adler2015pigz, ) of gzip, a commonly-used general compressor; (ii(N)Spr: Spring (chandak2018spring, ) and NanoSpring (Meng2023, ), state-of-the-art compressors for short and long reads, respectively; (iii(N)SprAC: (N)Spr integrated with a BWT-accelerator. There are various accelerators for BWT (e.g., (qiao2019fpga, ; guo2013gpu, ; zhao2017streaming, )). We consider an idealized accelerator that can fully eliminate the BWT execution time from (N)Spr; (iv0TimeDec: an idealized decompressor (with zero time needed for decompression, and with a compression ratio similar to the state-of-the-art genomic compressor), but inefficient for integration in resource-constrained environments (e.g., in the case of our evaluations, inside the SSD, for integration with the NDP genome analysis system); [^12] As a strongly conservative evaluation, 0TimeDec can also serve as an idealized representation of the closed-source industry software, called ORA (dragenora, ), recently developed for Illumina (illumina, ) short read (de)compression. ORA works similarly to other genomic compressors (as discussed in §2.2), but uses a different backend compressor for mismatch information, which still suffers from limitations for implementation in resource-constrained environments (detailed in §3.3). We cannot directly use ORA as our short-read decompression baseline because it decompresses both DNA bases and quality scores, whereas SAGe, like many other tools, targets only the bases (as discussed in §5.1.5). Since ORA’s source code and implementation details are unavailable, modifying it to target only bases is infeasible, which prevents a fair quantitative comparison against it. (vSGSW: SAGe with its decompression in software, running on the host system, to show the benefits of SAGe’s algorithmic optimizations; and (viSG: SAGe’s full implementation, with its decompression implemented in hardware; (viiSGin: SG with its hardware implemented in the SSD, to integrate with the NDP genome analysis system on the same chip. [^13] Other configurations connect to the genomic accelerator using PCIe interface.
@@
Datasets. We use real-world genomic data, with both short and long read sets. Storage systems must handle numerous read sets (e.g., in a medical center (Bick2024, ; Li2023whole, ) or samples in a cohort (Bick2024, ; Li2023whole, ; 10002015global, ; uk10k2015uk10k, )), so compressing each read set is essential to reduce the overall burden, regardless of their individual sizes. Thus, we analyze read sets of varying sizes, as listed in Table 3.
@@ 7 GB/s sequential-read BW
Performance. We design a simulator that models all components involved during SAGe’s execution, including accessing storage, hardware components (both for SAGe’s hardware and genomic accelerators (chen2023gem, ; mansouri2022genstore, )), the SSD’s internal DRAM (for the NDP accelerator (mansouri2022genstore, )), host operations, and host/SSD/accelerators interfaces. Using this methodology, as also demonstrated in prior works (e.g., (mansouri2022genstore, ; park2022flash, ; megis, )), enables us to flexibly incorporate state-of-the-art system configurations in our analysis. We feed the latency and throughput of each component to this simulator. For the components in the hardware-based operations, we implement SAGe’s logic components in Verilog and synthesize them using the Synopsys Design Compiler (synopsysdc, ) with a 22 nm library (22gf, ). We use two state-of-the-art simulators, Ramulator (kim2016ramulator, ; ramulatorsource, ) to model the SSD’s internal DRAM, and MQSim (tavakkol2018mqsim, ; mqsimsource, ) to model the SSD’s internal operations. For the hardware read mapper, we use the throughput numbers reported by the original paper (chen2023gem, ). For the software-based operations (software decompressors), we measure performance with their best-performing thread counts, on a high-end real system, an ${}^{\text{\textregistered}}$ ${}^{\text{\textregistered}}$ 7742 CPU (amdepyc, ) with 128 physical cores and 1.5-TB DRAM.
SSD Configurations. In our real system experiments, we use a state-of-the-art performance-optimized SSD (samsungPM1735, ) with PCIe Gen4 (PCIE4, ) and a cost-optimized SSD (samsung870evo, ) with a SATA3 interface (SATA, ). For the MQSim simulations of the in-storage NDP tasks, we faithfully model the SSDs based on the configurations in Table 1.
@@ 8.1. Performance
Fig. 13 shows end-to-end performance. All configurations are integrated with GEM (chen2023gem, ) for read mapping. For SGin, we evaluate an additional configuration integrated with the NDP in-storage filter (ISF), GenStore (mansouri2022genstore, ). We use GenStore-EM for filtering exactly-matching short reads, and GenStore-NM for filtering non-matching long reads in a contamination search read mapping use case, presented in the original work (mansouri2022genstore, ). Speedup is normalized to the genomics-specific baseline (N)Spr. SGSW, SG, and SGin+ISF enable an ablation analysis of SAGe’s optimizations.
@@
We make five key observations. First, SAGe provides significant performance benefits. On the system with the PCIe (SATA) SSD, SG leads to 12.3$\times$ (8.1$\times$), 3.9$\times$ (2.7$\times$), and 3.0$\times$ (2.1$\times$) average speedup over pigz, (N)Spr, and (N)SprAC, respectively. Second, SG matches 0TimeDec in performance because SG fully hides the decompression overhead in the execution pipeline. As explained in §3, I/O operations, decompression, and read mapping are pipelined, so total throughput depends on the slowest stage. 0TimeDec and SG achieving the same performance shows that decompression is no longer the slowest stage of the end-to-end pipeline. Third, due to its efficient access patterns, SAGe’s implementation in software (SGSW) also leads to speedups; 2.3$\times$ on average over (N)Spr. However, SGSW’s decompression still significantly impacts end-to-end performance. SGSW leads to up to 4.0$\times$ slowdown compared to 0TimeDec and SG. The hardware implementation of SAGe decompression provides greater speedup than its software by handling bitwise operations (on arrays and guide arrays) more efficiently and managing data flow more effectively in a fine-grained manner.
Fourth, by efficiently integrating with the in-storage filter, SGin+ISF leads to 7.8$\times$ (2.5$\times$) average speedups compared to (N)SprAC on the system with the PCIe (SATA) SSD. SGin+ISF leads to better performance than SG in all cases, except when both of the following conditions are met: (i) the input and application do not largely take advantage of in-storage NDP (i.e., in this case, ISF does not filter many reads in the read set), and (iithe SSD’s limited external bandwidth bottlenecks performance (e.g., RS1 and RS4 with the SATA SSD). In these cases, the SG configuration should be used for better performance by decompressing data outside the SSD, thus avoiding moving larger decompressed data through the limited-bandwidth interface. Fifth, regardless of how much decompression tools are optimized for performance, if they are unsuitable for adoption in resource-constrained environments (as discussed in §3.3), they miss out on the benefits of a wide range of genome analysis systems. For example, even 0TimeDec leads to on average 1.8$\times$ (up to 9.9$\times$) lower performance than SG+ISF.
Data Preparation. Fig. 14 shows only data preparation throughput (with the PCIe SSD), normalized to pigz. SG leads to 91.3$\times$, 29.5$\times$, and 22.3$\times$ average speedups over pigz, (N)Spr, and (N)SprAC, respectively. As shown in Fig. 13, this faster data preparation leads to large improvements in end-to-end genome analysis throughput.
Figure 14. Data preparation speedup.
@@
Performance with Multiple SSDs. Since all data structures in SAGe are accessed sequentially, they can be disjointly partitioned between multiple SSDs. Fig. 15 shows the performance of SG and SGin+ISF with multiple PCIe SSDs, normalized to (N)Spr. First, SG maintains its large speedup when data is partitioned among multiple SSDs. Second, SGin+ISF’s benefits improve for some datasets (RS3 and RS5) since more SSDs improve the performance of ISF, which is at the critical path of their end-to-end execution. We conclude that through efficient partitioning, SAGe is able to effectively leverage multiple SSDs. Leveraging this partitioning, and given that each genomic read can be analyzed independently, SAGe’s design is also compatible for adoption in distributed storage systems.